OpenStudy (aonz):

Use the discriminant to find the gradients of the lines that pass through the point (7, 1) and are tangent to the circle x^2 + y^2 = 25.

OpenStudy (aonz):

@satellite73 help please ^^

OpenStudy (anonymous):

ok, we need \(y'\) using implicit diff you get \[2x+2yy'=0\] solve for \(y'\) gives \(y'=-\frac{x}{y}\) and now plug in the numbers

OpenStudy (anonymous):

you could also solve for \(y\) in this example, and get \[y=\pm\sqrt{25-x^2}\] then take the derivative, then plug in the numbers try it you will see you get the same answer

OpenStudy (aonz):

y' = dy/dx right?

OpenStudy (anonymous):

and the third method is to recognize this as a circle and know that the answer is always \(-\frac{x}{y}\) for a circle

OpenStudy (anonymous):

yeah, \(y'\) takes two keystrokeds, \(\frac{dy}{dx}\) takes 12

OpenStudy (aonz):

ok the question asks to use the discriminant.. which way would that be?

OpenStudy (anonymous):

i have absolutely no idea the only definition i know for "discriminant" is \(ax^2+bx+c=0\) then the discriminant is \(b^2-4ac\)

OpenStudy (anonymous):

you are being asked for the slope of the line tangent to the circle idea is to find the derivative, plug in the numbers

OpenStudy (aonz):

(7, 1) and a tangent to the circle x^2 + y^2 = 25 Is it possible to find the equation of this line?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you have a point, so all you need is the slope

OpenStudy (anonymous):

is it clear how to find the slope, or no?

OpenStudy (aonz):

yep all clear

OpenStudy (aonz):

slope is 25/28?

OpenStudy (anonymous):

ooh lets back up

OpenStudy (anonymous):

oh crap i read the problem wrong my fault \((7,1)\) is not a point on the circle is it? you want the equation for the line tangent to the circle through the point \((7,1)\)

OpenStudy (aonz):

yea i think so

OpenStudy (anonymous):

a point on the circle will look like \[(x, \sqrt{25-x^2})\] and the slope will be \[\frac{\sqrt{25-x^2}-1}{x-7}\]

OpenStudy (anonymous):

this in turn must be equal to the derivative, which is \[-\frac{x}{\sqrt{25-x^2}}\]

OpenStudy (anonymous):

set them equal and solve for \(x)

OpenStudy (aonz):

hmm?

OpenStudy (anonymous):

you have a point on the circle it looks like \((x,\sqrt{25-x^2})\) or \((x,-\sqrt{25-x^2})\)

OpenStudy (anonymous):

then you have a point not on the circle, \((7,1)\)

OpenStudy (aonz):

ok i get that

OpenStudy (anonymous):

the slope between those two points is going to be \[\frac{\sqrt{25-x^2}-1}{x-7}\]

OpenStudy (aonz):

ok

OpenStudy (anonymous):

maybe it would be clearer if i replaced \(x\) by \(a\) but lets keep with the \(x\)

OpenStudy (anonymous):

now we take the derivative of \[y=\sqrt{25-x^2}\]

OpenStudy (anonymous):

and get \[y'=\frac{-x}{\sqrt{25-x^2}}\]

OpenStudy (anonymous):

ok so far?

OpenStudy (aonz):

yep

OpenStudy (anonymous):

now you want the line through \((7,1)\) tangent to the circle at some point if it is tangent to the circle, then the slope must be equal to the derivative at the point where it touches the circle

OpenStudy (aonz):

im pretty sure there would be a quicker way because they rest of the question are pretty basic :/ and not this complex lol

OpenStudy (anonymous):

in other words \(\frac{\sqrt{25-x^2}-1}{x-7}\) must be equal to \(\frac{-x}{\sqrt{25-x^2}}\)

OpenStudy (anonymous):

yeah i am thinking the same thing, but i don't see it

OpenStudy (anonymous):

maybe solve two equation \[x^2+y^2=25\] \[\frac{y-1}{x-7}=-\frac{x}{y}\]

OpenStudy (anonymous):

yeah, that might be easier !!

OpenStudy (aonz):

here is the worked example. i think its similiar

OpenStudy (anonymous):

does the solution look similar to the two equations i wrote above?

OpenStudy (anonymous):

yeah same idea

OpenStudy (anonymous):

by the way these equations are actually identical to my complicated method, because \(y=\sqrt{25-x^2}\) but i can see that this will be much easier

OpenStudy (aonz):

ok i got a answer! ^^ Thank you so much for your time and your effort!!

OpenStudy (anonymous):

lol what is it?

OpenStudy (aonz):

i thought i got it but i ended up with a negative square root :(

OpenStudy (aonz):

so i did y-1 = m(x-7) first right?

OpenStudy (anonymous):

ooh i see learn something new every day! in order for the line to be tangent, it means it must touch therefore the discriminant must be zero!! now i see what you were saying at the beginning!!!

OpenStudy (aonz):

then i got y = mx-7m+1

OpenStudy (anonymous):

yeah i am there too

OpenStudy (aonz):

now do i have to square that? cause that will be messy

OpenStudy (anonymous):

now i am at the step where you set \[(mx-7m+1)^2+x^2=25\]

OpenStudy (aonz):

well im going to expand it and see what i get :/ this is annoying

OpenStudy (anonymous):

yeah i am going to do it as well

OpenStudy (anonymous):

\[m^2 x^2-14 m^2 x+49 m^2+2 m x-14 m+1+x^2=25\]

OpenStudy (aonz):

hehe, im gonna double check my answers with wolfram ^^

OpenStudy (anonymous):

double check? i just used wolfram!

OpenStudy (anonymous):

now lets get the quadratic equation in \(x\)

OpenStudy (aonz):

lol... i did it on paper first :P

OpenStudy (anonymous):

\[(m^2+1)x^2+(-14m^2+2m)x+(49m^2-14m-24)=0\] really??

OpenStudy (anonymous):

i thought this was supposed to be the easy way!!

OpenStudy (anonymous):

now we set \(b^2-4ac=0\) right?

OpenStudy (aonz):

yea... wow... this is no easy way

OpenStudy (anonymous):

\((-14m^2+2m)^2-4(m^2+1)(49m^2-14m-24)=0\)

OpenStudy (anonymous):

i think i messed up somewhere i bet it is actually easier to use the derivative

OpenStudy (aonz):

Maybe their is a easier way using the discriminant...

OpenStudy (aonz):

anyways ill ask my teacher tomorrow and see what she says... Thanks for your help!!!

OpenStudy (anonymous):

sorry i will think about it and get back to you

OpenStudy (aonz):

If if you dont have time, dont worry about it :)

OpenStudy (aonz):

I got to sleep its 1:25 in the morning

OpenStudy (anonymous):

ok if you see this when you get up, it is not that hard at all solve \[x^2+y^2=25\\ \frac{y-1}{x-7}=-\frac{x}{y}\] and get \[x=3,y=4\] or \[x=4,y=-3\]